Test the hypothesis using the six step process.
Ho: p = 0.9 versus H1: p does not equal 0.9
n = 500, x = 440, a =0.05

Answer by stanbon(75887)

Test the hypothesis using the six step process.
Ho: p = 0.9 versus
H1: p does not equal 0.9
-------------
n = 500, x = 440, a =0.05
--------
p-hat = 440/500 = 0.88
---
z(0.88) = (0.90-0.88)/sqrt[0.9*0.1/500] = -1.4907
---
p-value = 2*P(z< -1.4907) = 0.1360
---
Since p-value is greater than 5%, fail to reject Ho.
===============
Cheers,
Stan H.
==============